3. Fifteen employees of a large manufacturing company were involved in a study to test a new production method. The mean production rate for the sample of 15 employees was 63 components per hour and the standard deviation was 8 components per hour. Calculate the margin of error for the mean number of components produced per hour at a 5% level of significance.


1
Expert's answer
2021-10-31T18:33:06-0400

Lets mark the margin of error as m, then:

"m = Cr*{\\frac {\u03c3\u200e} {\\sqrt{n}}}" , where Cr - critical value, σ‎ - standart deviation of the sample, n - sample size

σ‎ is known, then we have to find Cr

Since we have small sample size(<30) and the population distribution is not known, it is appropiate to use t-statistics with n-1 = 14 degrees of freedom

It is also appropriate to make a two-sided test, so

"P(""t(14)>Cr) = {\frac a 2}=0.025" , a - level os significance(0.05)

The t-values can be found from the tables

Cr = 2.145

"m= 2.145*{\frac 8 {\sqrt{15}}}=0.107=4.43"

That means that with 95% confidence the mean value lies in the interval (63-4.43, 63+4.43) =

= (58.57, 67.43)